OpenStudy (shal):

Can someone help me with 1 math question? Plzz...Ill give medal...

OpenStudy (shal):

If a line segment has endpoints (2, –5) and (–1, –1), what is its slope? (Points : 1) –3/4 –4/3 –1/6 –6/1

OpenStudy (anonymous):

m=(-5-(-1))/(2-(-1))=-4/3

OpenStudy (shal):

If a line segment has endpoints (2,-5) and (-1,-1), What is its slope? a.) b.) c.) or d.)

OpenStudy (shal):

@AJ01

OpenStudy (anonymous):

B

OpenStudy (shal):

are yu sure?

OpenStudy (shal):

Can yu help me with 2 more?

OpenStudy (shal):

A line has a slope of 2 and passes through the point (4, 0). What are the coordinates of another point that is on the line? (Points : 1) (2, 0) (0, 2) (4, 6) (6, 4)

OpenStudy (anonymous):

yes.... slope=(y-y1)/(x-x1)=(-5-(-1))/(2-(-1))=(-5+1)/(2+1)=-4/3

OpenStudy (anonymous):

okay

OpenStudy (shal):

aw Thanks!

OpenStudy (shal):

so how bout the other one?

OpenStudy (shal):

@AJ01

OpenStudy (anonymous):

findinf line eq. y-y1=m(x-x) y-(0)=2(x-(4) y=2x-8 D is the answer if you put x=6 in the above eq you will get 4

OpenStudy (shal):

can u help me with the last one? Plzzzzzzzzzzzzzzzzzz

OpenStudy (anonymous):

ok

OpenStudy (shal):

THX!

OpenStudy (shal):

What is the slope of the line that passes through the points (1, 2) and (-2, 4)? (Points : 1) -2/3 -3/2 2 –2

OpenStudy (shal):

d.) is -2 sorry it came out so bad

OpenStudy (anonymous):

okay slope =(2-4)/(1-(-2))=-2/3

OpenStudy (shal):

Thank YOU SO MUCH! ;)

OpenStudy (anonymous):

:)

OpenStudy (shal):

i gave uu a medal

OpenStudy (shal):

and became a fan of u

OpenStudy (anonymous):

THX

OpenStudy (shal):

np :)

OpenStudy (anonymous):

THX again

OpenStudy (shal):

OH WAIT!!!!!!!!!!!!! CAN YU HELP ME WITH 2 MORE I GOT THEM WRONG!!!!!!!

OpenStudy (shal):

@AJ01

OpenStudy (shal):

Can ANYONE HELP ME WITH GEOMETRY QUESTIONS?!

OpenStudy (anonymous):

ok

OpenStudy (shal):

Oh thx

OpenStudy (shal):

Thank yu so much i dont know what i would do....I got these two questions wrong...

OpenStudy (anonymous):

what are they

OpenStudy (shal):

Therr in the doc, geo questions

OpenStudy (shal):

2 questions

OpenStudy (shal):

so for the first one, i got 1/4 which i got wrong and the second one i got 5/4, which was wrong...

OpenStudy (anonymous):

first one asking about the line EF okay

OpenStudy (shal):

yeh

OpenStudy (anonymous):

reading the points from the plot

OpenStudy (shal):

ok

OpenStudy (anonymous):

what E should be? and F?

OpenStudy (shal):

4/5?

OpenStudy (anonymous):

i am not asking about the slope i am asking about the points

OpenStudy (shal):

oh ok

OpenStudy (shal):

F is 2,-2 and E is -6,8 right?

OpenStudy (anonymous):

that is right slope= (y-y1)/(x-x1)

OpenStudy (shal):

yes

OpenStudy (anonymous):

slope=(-2-8)/(2-(-6))=-10/4=-5/4

OpenStudy (anonymous):

do the other one

OpenStudy (shal):

oh ok check if its right ok?

OpenStudy (shal):

for the next one

OpenStudy (anonymous):

i am waiting :)

OpenStudy (shal):

ok :)

OpenStudy (shal):

-1/4?

OpenStudy (anonymous):

what are the points C and D

OpenStudy (shal):

c is 5.-2 and d is 7,6

OpenStudy (anonymous):

that is right

OpenStudy (shal):

ok,yay

OpenStudy (shal):

and then?

OpenStudy (anonymous):

points are okay but not the slope right the slope

OpenStudy (shal):

4/1?

OpenStudy (anonymous):

yes that is right

OpenStudy (shal):

YAAY!

OpenStudy (shal):

THNKS! FOR THE HELP!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! I APPRECIATE YOUR TIME AND EFFORT IN HELPING ME! :)

OpenStudy (anonymous):

no problem......... take care

OpenStudy (shal):

you too